OpenStudy (anonymous):

A circle has a diameter of twelve millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth. thirty-seven point seven millimeters; four hundred fifty-two point two square millimeters eighteen point eight millimeters; one hundred thirteen point zero square millimeters thirty-seven point seven millimeters; one hundred thirteen point zero square millimeters eighteen point eight millimeters; four hundred fifty-two point two square millimeters

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

It makes it a real pain to have the numbers written out in plain t ext. All the same, the formulas are: \[\text{Area=}\pi r^2 \] \[\text{Circumference=}2\pi r\]

OpenStudy (anonymous):

thans i get it now